OpenStudy (anonymous):

Using this f(x) = x2 + 6x – 16 This is algebra 2 I get how to find the vertex once its put in the right form but i dont know how to put it in this form f(x) = a(x - h)2 + k

OpenStudy (phi):

complete the square. Have you learned how?

OpenStudy (anonymous):

yes i have i have to get my notes real quick

OpenStudy (anonymous):

okay i got my notes but im not sure what to do with it i think what you would do is get rid of the coefficient

OpenStudy (phi):

if you have time, this goes into details https://www.khanacademy.org/math/algebra/quadratics/completing_the_square/v/solving-quadratic-equations-by-completing-the-square meanwhile, if you have f(x) = x^2 + bx + c (and you do here) find the b (number in front of the x) , and divide it by 2 then square it. (b/2)^2 then put in f(x) = x^2 + bx + (b/2)^2 - (b/2)^2 + c

OpenStudy (phi):

the first three terms in f(x) = x^2 + bx + (b/2)^2 - (b/2)^2 + c i.e. the x^2 + bx + (b/2)^2 can be re-written as (x+b/2)^2

OpenStudy (anonymous):

f(x) = x2 + 6x – 16 so 6 is the b and c is the -16 ?

OpenStudy (anonymous):

that was supposed to come out a -16 but it didn't show up right on my screen

OpenStudy (phi):

OS is a bit screwed up. It shows up if you re-fresh the browser (which seems to take forever) meanwhile, yes b is 6 you now need b/2

OpenStudy (anonymous):

so 3

OpenStudy (phi):

now square it

OpenStudy (anonymous):

9

OpenStudy (phi):

f(x) = x^2 + 6x – 16 now add in +9 - 9 (which is zero, so it's ok) into the equation f(x) = x^2 + 6x +9 - 9 – 16 and look at only the first 3 terms f(x) = (x^2 + 6x +9) - 9 – 16

OpenStudy (phi):

refresh your browser (reload this page) if you can't read that

OpenStudy (anonymous):

why would i need do what i just did though ? like with the 6/2 and squaring that

OpenStudy (anonymous):

and why wouldn't it just cancel out ?

OpenStudy (phi):

I posted a link to a Khan video that explains why this works. meanwhile, we are just doing it.

OpenStudy (anonymous):

wouldn't you add 9 to both sides thats what it showed in my example

OpenStudy (phi):

we could cancel it out. but we don't have to. Instead , we will group the first 3 terms and re-write them f(x) = (x^2 + 6x + 9) - 9 -16 the -9 - 16 can be simplified to -25 f(x) = (x^2 + 6x + 9) -25 and because of adding in +9, the first part is a perfect square, and can be written as (x+ 3)^2 notice: (x+3)(x+3)= x^2 +3x +3x + 9 = x^2 +6x + 9

OpenStudy (anonymous):

im not getting this at all ill add a picture of what my example says

OpenStudy (phi):

yes we could add +9 to both sides which makes more sense if we had x^2 + 6x -16 = 0 and we did x^2 + 6x+9 -16 = 0+9 but notice if we now add - 9 to both sides it would look like this x^2 + 6x+9 -9 -16 = 0+9-9 or x^2 + 6x+9 -9 -16 = 0 which is ok.

OpenStudy (anonymous):

thats what i got it from

OpenStudy (phi):

yes, that is fine. if you like we can do that: f(x) = x^2 + 6x -16 add +9 to both sides (9 is from (6/2)^2 i.e. (b/2)^2 ) f(x)+9 = x^2 + 6x +9 -16 if we want to "get rid of the -16" we can add +16 to both sides f(x)+9+16= x^2 + 6x +9 -16 + 16 and simplify on the left 9+16 is 25 on the right 16-16 is 0 we get f(x) +25 = x^2 + 6x +9

OpenStudy (phi):

ok ?

OpenStudy (anonymous):

and the 9 was from the 6/2 squared ?

OpenStudy (phi):

yes

OpenStudy (anonymous):

okay i got that

OpenStudy (phi):

the important idea is if you have y = x^2 + bx + other junk if we add in (b/2)^2 (and subtract it somewhere's else to keep things balanced) like this y + (b/2)^2 = x^2 + bx + (b/2)^2 + other junk this part: x^2 + bx + (b/2)^2 can be written as (x+b/2)^2 y + (b/2)^2 = (x+b/2)^2 + other junk

OpenStudy (phi):

if that is not clear, when you have time, watch the video. meanwhile, you should have for your problem f(x) +25 = x^2 + 6x +9 f(x) + 25 = (x+3)^2 and add -25 on both sides to "clean up" f(x) +25 - 25 = (x+3)^2 - 25 f(x) = (x+3)^2 - 25

OpenStudy (anonymous):

so that was it or is there more ?

OpenStudy (anonymous):

f(x) = a(x - h)2 + k in this form it would be f(x)=(x-3)^2 -25 but i dont see where a is at if thats what that equation is for

OpenStudy (phi):

if you need the vertex, you can match up that equation with f(x) = a(x-h)^2 + k vertex is (h,k)

OpenStudy (anonymous):

did i do it right up there

OpenStudy (anonymous):

h would be -3 and k would be -25 ?

OpenStudy (anonymous):

I get what you did but i still dont see where the letters match up at

OpenStudy (phi):

(x) = (x+3)^2 - 25 to make it "match up" with f(x) = a(x - h)2 + k we write f(x) = (x+3)^2 - 25 as f(x) = 1*(x- (-3) )^2 + -25 match to a*(x- h)^2 + k

OpenStudy (phi):

a matches 1 h matches (-3) or just -3 k matches -25

OpenStudy (anonymous):

okay so the vertex would be (-3,-25) right ?

OpenStudy (phi):

yes

OpenStudy (anonymous):

and since a didnt show up that just got replaced with a 1?

OpenStudy (phi):

I would say we don't bother to write down the 1*stuff in other words if we have 1*x we would just say x rather than 1x so if you see (x- -3)^2 it's the same as 1*(x- -3)^2

OpenStudy (anonymous):

would a always be like that or could it have a higher number than one

OpenStudy (phi):

"a" will make the parabola fatter (or skinnier) than normal if it's not 1

OpenStudy (phi):

we would get a not 1 if we started with for example 2x^2 +6 x -16 but it makes "completing the square" harder. (so you don't want to see a not be 1!)

OpenStudy (anonymous):

so a in that would be the 2

OpenStudy (phi):

yes, in that example, a would end up being 2 btw, if you have f(x) = a x^2 + bx +c the x value of the vertex is -b/(2a) for your problem we get -6/2= -3 (much faster than completing the square) to find the y value of the vertex, we would set x= -3 in the equation and solve for f(-3)

OpenStudy (anonymous):

so to summarize how to figure it out i would say what i did was take 6x and divided it by 2 then squared the answer to get 9 and with to get rid of -16 i added +16 on both sides and got f(x) 25 = x^2 (x+3)^2 the to clean it up i subtracted 25 from both sides and ended up with f(x)=(x+3)^2-25

OpenStudy (phi):

take the 6 from 6x (just to be picky) divide by 2 to get 3 square to get 9 next we add 9 to the right-side of the equation. to keep things balanced, we *either* add 9 to the left-side OR add -9 also to the right-side

OpenStudy (anonymous):

Sorry it kicked me out so i had to go on my other account